# Image Classification in High-Energy Physics: A Comprehensive Survey of Applications to Jet Analysis

### v2

@@ 3 HEP Jet image representation
This section provides an overview of the Jet datasets comprising various forms of Jet images obtained and generated through different methods. Additionally, the current section delves into diverse pre-processing and feature extraction techniques employed in this context.
3.1 Available datasets
@@
The conseil Européen pour la recherche nucléaire (CERN) open data portal provides access to a variety of datasets from experiments conducted at the Large LHC. These datasets include information about collisions, particles, and Jet images. The portal offers a great starting point for those interested in HEP datasets. Figure 3 illustrates samples of Jet images, featuring the average of $p_{T}$-normalized quark and gluon Jet images across 5 distinct $\chi$ bins. The Jet images may undergo different preprocessing techniques, discussed later, prior to input into ML/DL models for classification or prediction tasks. Table LABEL:tab:datasets showcases the datasets most commonly utilized in the reviewed research within this paper, including some generated datasets.generated.
@@
Table 3: A summary of available data set and EG for Jet HEP.
Dataset/EG Name Description DLA ?
@@
Dataset ATLAS open data Is one of the largest particle physics experiments at the LHC. They offer an "Open Data" initiative with datasets that include collision data and simulated samples. These datasets can be used to study Jet images and other particle physics phenomena. Yes [^1] http://opendata.cern.ch/search?page=1&size=20&experiment=ATLAS
Dataset CMS open data Compact muon solenoid (CMS) is another major experiment at the LHC. Similar to ATLAS, CMS provides open data for educational and research purposes. The datasets include information about collisions, particles, and Jets. Yes  [^2] https://opendata.cern.ch/search?page=1&size=20&q=jet%20images&experiment=CMS
EG Delphes simulation Is a particle physics event generator designed to produce simulated collision events that are similar to those observed in real experiments. It includes tools to generate Jet images based on the data produced in simulations. Yes [^3] https://cp3.irmp.ucl.ac.be/projects/delphes
EG MadGraph Is a popular event generator used in particle physics simulations. It can generate events involving Jets and other particles, which can then be turned into Jet images. Yes [^4] http://madgraph.phys.ucl.ac.be/
EG FASTSim Is a tool for simulating high-energy particle collisions. It can generate Jet images from simulated collision events and is often used for studying ML techniques in HEP. Yes [^5] https://twiki.cern.ch/twiki/bin/view/CMSPublic/SWGuideFastSimulation
@@
Dataset Complete It belongs to CERN and contains muon, kaon, pion, electron, and proton. In the complete dataset training, 400,000 Jets are used for training, and the remaining 290,000 are used for testing and assessing performance [25]. No
Dataset Monte Carlo It is generated through a dependable framework, created by integrating various tools like Pythia 8 for generating HEP events, Delphes for emulating the detector’s response, and RAVE for reconstructing secondary vertices [29]. No
@@
Dataset Top tagging This dataset comprises 1.2 million training samples, 400,000 for validation, and another 400,000 for testing. Each entry in this dataset corresponds to an individual Jet, with its source being either an energetic top quark, a light quark, or a gluon. These events were generated using the PYTHIA8 Monte Carlo event generator, and the response of the ATLAS detector is simulated using the DELPHES software package Yes [^6] https://zenodo.org/record/2603256
@@
Dataset Quark-gluon tagging The dataset is created by generating signal (quark) and background (gluon) Jets through PYTHIA8. For the signal Jets, the process involves $Z(\rightarrow νν)+(u,d,s)$, and for the background Jets, it uses $Z(\rightarrow νν)+g$. Notably, there is no simulation of the detector. The particles that are not neutrinos in the final state are grouped into Jets using the anti-kT algorithm with a radius parameter of $R=0.4$. In total, this dataset contains 2 million Jets, evenly split between signal and background categories [30]. No
@@
Dataset Higgs dataset The dataset originates from Monte Carlo simulations. The initial 21 attributes (found in columns 2-22) represent particle detector-derived kinematic properties within the accelerator. The remaining seven attributes are transformations of the initial 21, constituting high-level features engineered by physicists to aid in distinguishing between the two categories. Yes [^7] https://archive.ics.uci.edu/dataset/280/higgs
@@
Dataset QCD multi-Jet Samples are generated across different ranges of scalar sum of $p_{T}$, namely 1000-1500 GeV, 1500-2000 GeV, and 2000-Inf GeV. After excluding samples with $p_{T}$ values less than 1000 GeV, the dataset consists of around 450. 103 training images, 150.103 validation images, and 150.103 testing images [16]. No
@@ 3.2 Pre-processing of Jet images
The subsequent sequence of data-driven preprocessing procedures was employed on the Jet images:
@@ 3.3 Feature extraction and selection
Di Luca et al. [29] presents an automated feature selection procedure for particle Jet classification in HEP experiments. The authors use ML boosted tree algorithms to rank the importance of observables and select the most important features associated with a particle Jet. They apply this method to the specific case of boosted Higgs boson decaying to two b-quarks ($H\rightarrow bb$) tagging and demonstrate the impact of feature selection on the performance of the classifier to distinguish these events amidst the substantial and unalterable background originating from quantum chromodynamics (QCD) multi-Jet production. They also train a fully connected neural network to tag the Jets and compare the results obtained using all the features or only those selected from the procedure which consists of two main steps: data preparation and feature ranking extraction. The authors discover that the azimuthal angles of the large-R Jet and the variable radius (VR)-track Jets appear towards the end of the feature ranking. At the top of the ranking, they find the $p_{T}$ of the two VR-track Jets, along with certain details regarding the secondary vertex, such as its mass, energy, and displacement. The study shows that selecting the highest-ranked features achieves performance nearly as effective as that of the full model, with only a slight deviation of a few percent. This approach can be expanded to accommodate the increased number of observable variables that upcoming collider experiments will gather from high $p_{T}$ particle Jets. The data for this research comes from proton-proton collision events featuring a boosted Higgs boson that decays into two $b$ quarks. In [33], solutions have been proposed for classifying events extracted from the 2014 Higgs ML Kaggle dataset [^8] https://www.kaggle.com/c/higgs-boson. The dataset includes a mix of low-level and high-level attributes: it contains 18 low-level features that include three-dimensional momenta ($p_{T}$, $\eta$, $\phi$), missing transverse momentum, and the total transverse momentum from all jets; additionally, there are 13 high-level features motivated by physics, covering invariant masses and angular separations among objects in the final state. Table LABEL:tab:features summarizes the features utilized, which hold potential for future application within the context of HEP. The authors aim to ensure that the suggested networks make effective use of low-level information; otherwise, there’s a risk of losing these features during selection. Their focus lies in determining the necessity of high-level features. The proposed DNN model effectively utilize the low-level information in the data and autonomously learn their own high-level representations. Boost-invariant polynomial (BIP) features are a type of mathematical representation used in HEP for analyzing particle collision data. They are constructed to be invariant under boosts, meaning they remain unchanged under transformations to different reference frames with different velocities. These features are designed to capture important characteristics of particle jets, such as their energy distribution and substructure, while ensuring consistency across various experimental conditions. BIP features are particularly useful for tasks like jet tagging and classification in HEP experiments, as employed in [34].
@@
4 Available AI models for HEP Jet images
Many DL architectures have been proposed in the SOTA of HEP domain to identify particles, Table 5 summarizes and compares the most efficient DL models, used in HEP, based on their architectures and performances. For example, the work in [35] proposed PartT, which is a new Transformer-based architecture for Jet tagging. Its main task is to identify the origin of a Jet of particles produced in HEP experiments. ParT makes use of two sets of inputs: (i) the particle input, which includes a list of features for every particle and forms an array, and (ii) the interaction input, which is a matrix of features for every pair of particles. ParT employs a novel pairwise multi-head attention (P-MHA) mechanism, which allows the model to attend to pairs of particles and learn their interactions. The P-MHA is more effective than standard plain multi-head attention. This assertion is substantiated when the pre-trained ParT models are fine-tuned on two widely adopted Jet tagging benchmarks, the quark-gluon tagging dataset and the binary classification dataset for identifying boosted $W$ bosons decaying to two quarks. The fine-tuning process involves training the ParT models on a smaller labeled dataset specific to each benchmark, which allows the models to learn the specific features and patterns relevant to each task. The fine-tuned ParT models achieve significantly higher tagging performance than the models trained from scratch and outperform the previous SOTA models, including ParticleNet and other Transformer-based models.
Another architecture called the LorentzNet is proposed in [30], which is based on the Lorentz group equivariant block (LGEB) block. The structure of LGEB consists of several layers, including Minkowski norm and inner product, sum pooling, a multi-Layer perceptron (MLP), and a Clebsch-Gordan tensor product. The input of LGEB is a set of 4-momentum vectors, which are transformed by the Minkowski norm and inner product layer to obtain Lorentz-invariant geometric quantities. The sum pooling layer aggregates the geometric quantities to obtain a scalar representation of the input. The MLP layer is used to learn a nonlinear mapping from the scalar representation to a new feature space. Finally, the Clebsch-Gordan tensor product layer is used to combine the new feature space with the original input to obtain the output of LGEB. It is designed as a Lorentz group-equivariant mapping to preserve the symmetries of the Lorentz group, ensuring the model’s equivariance and universality.
Unlike the previous models, the authors in [36] proposed the equivariant graph neural networks (EGNN) model, which is a graph neural network (GNN) architecture that is translation, rotation, and reflection equivariant (E(n)), and permutation equivariant with respect to an input set of points. It uses a set of filters that are equivariant to the action of the symmetry group, which are constructed using a combination of radial basis functions and Chebyshev polynomials. The EGNN algorithm possesses the same flexibility as the GNN technique, while also maintaining E(n) equivariance similar to the radial field algorithm. Additionally, it eliminates the requirement for computationally intensive procedures, such as spherical harmonics. The EGNN exceeds other equivariant and non-equivariant options while maintaining efficiency in terms of running time. Moreover, the EGNN approach demonstrates a 32% reduction in error compared to the SOTA method.
Figure 6: The architecture of EGNN model suggested in [36].
Similarly, [37] proposed the ParticleNet model. The architecture is a customized neural network that operates directly on particle clouds for Jet tagging. It uses dynamic graph convolutional neural networks (CNNs) to process the unordered set of constituent particles that make up a Jet. The architecture consists of three EdgeConv blocks, each with a different number of channels and nearest neighbors. EdgeConv block starts by representing a point cloud as a graph, whose vertices are the points themselves, and the edges are constructed as connections between each point to its K-nearest neighbors (KNN) points. The EdgeConv block then finds the KNN particles for each particle, using the "coordinates" input of the EdgeConv block to compute the distances. Inputs to the EdgeConv operation, the "edge features," are constructed from the "features" input using the indices of KNN particles. The EdgeConv procedure is executed using a three-layer MLP. Each layer is structured to include a linear transformation, succeeded by batch normalization, and subsequently a rectified linear unit (ReLU) activation. Additionally, a shortcut connection is integrated into every block parallel to the EdgeConv operation, facilitating the direct passage of input features. An EdgeConv block is defined by two key hyper-parameters: the neighbor count $k$ and the channel count $C$, which respectively denote the number of neighbors to consider and the number of units within each layer of linear transformation. The EdgeConv blocks play a crucial role in learning the local features of the particle cloud and aggregating them into a global feature vector for the Jet. Following EdgeConv blocks, global average pooling aggregates particle features, leading to a 256-unit fully connected layer, ReLU activation, dropout, and a 2-unit softmax output for binary classification. The ParticleNet architecture achieves SOTA performance on two representative Jet tagging benchmarks and is improved significantly over existing methods.
Figure 7: The architecture of ParticleNet model suggested in [37].

### v3

@@ 3 HEP Jet representation
This section provides an overview of the Jet datasets comprising various forms of Jet data obtained and generated through different methods. Additionally, the current section delves into diverse pre-processing and feature extraction techniques employed in this context.
3.1 Available datasets and simulation tools
@@
The conseil Européen pour la recherche nucléaire (CERN) open data portal provides access to a variety of datasets from experiments conducted at the Large LHC. These datasets include information about collisions, particles, and Jet images and PCs. The portal offers a great starting point for those interested in HEP datasets. Figure 3 illustrates samples of Jet images, featuring the average of $p_{T}$-normalized quark and gluon Jet images across 5 distinct $\chi$ bins. The Jet images or PC may undergo different preprocessing techniques, discussed later, prior to input into ML/DL models for classification or prediction tasks. Table LABEL:tab:datasets presents the datasets, along with several simulation tools, most commonly used in the research reviewed in this paper.
@@
Table 3: A summary of available datasets, and simulation tools for Jet HEP analysis.
Name Description DLA ?
@@
Datasets ATLAS open data Is one of the largest particle physics experiments at the LHC. They offer an "Open Data" initiative with datasets that include collision data and simulated samples. These datasets can be used to study Jet images and other particle physics phenomena. Yes [^1] http://opendata.cern.ch/search?page=1&size=20&experiment=ATLAS
[…]
@@
Complete It belongs to CERN and contains muon, kaon, pion, electron, and proton. In the complete dataset training, 400,000 Jets are used for training, and the remaining 290,000 are used for testing and assessing performance [26]. No
@@
Top tagging This dataset comprises 1.2 million training samples, 400,000 for validation, and another 400,000 for testing. Each entry in this dataset corresponds to an individual Jet, with its source being either an energetic top quark, a light quark, or a gluon. These events were generated using the PYTHIA8 Monte Carlo event generator, and the response of the ATLAS detector is simulated using the DELPHES software package Yes [^3] https://zenodo.org/record/2603256
@@
Quark-gluon tagging The dataset is created by generating signal (quark) and background (gluon) Jets through PYTHIA8. For the signal Jets, the process involves $Z(\rightarrow νν)+(u,d,s)$, and for the background Jets, it uses $Z(\rightarrow νν)+g$. Notably, there is no simulation of the detector. The particles that are not neutrinos in the final state are grouped into Jets using the anti-kT algorithm with a radius parameter of $R=0.4$. In total, this dataset contains 2 million Jets, evenly split between signal and background categories [31]. No
@@
Higgs dataset The dataset originates from Monte Carlo simulations. The initial 21 attributes (found in columns 2-22) represent particle detector-derived kinematic properties within the accelerator. The remaining seven attributes are transformations of the initial 21, constituting high-level features engineered by physicists to aid in distinguishing between the two categories. Yes [^4] https://archive.ics.uci.edu/dataset/280/higgs
@@
QCD multi-Jet Samples are generated across different ranges of scalar sum of $p_{T}$, namely 1000-1500 GeV, 1500-2000 GeV, and 2000-Inf GeV. After excluding samples with $p_{T}$ values less than 1000 GeV, the dataset consists of around 450. 103 training images, 150.103 validation images, and 150.103 testing images [17]. No
Simulation tools Delphes Is a particle physics event generator designed to produce simulated collision events that are similar to those observed in real experiments. It includes tools to generate Jet based on the data produced in simulations. Yes [^5] https://cp3.irmp.ucl.ac.be/projects/delphes
MadGraph Is a popular event generator used in particle physics simulations. It can generate events involving Jets and other particles, which can then be turned into Jet PC or images. Yes [^6] http://madgraph.phys.ucl.ac.be/
FASTSim Is a tool for simulating high-energy particle collisions. It can generate Jets from simulated collision events and is often used for studying ML techniques in HEP. Yes [^7] https://twiki.cern.ch/twiki/bin/view/CMSPublic/SWGuideFastSimulation
Monte Carlo It is generated through a dependable framework, created by integrating various tools like Pythia 8 for generating HEP events, Delphes for emulating the detector’s response, and RAVE for reconstructing secondary vertices [32]. No
@@ 3.2 Pre-processing for ML-based Jet analysis
The subsequent sequence of data-driven preprocessing procedures was employed on the Jet images and can also be adapted for PCs):
@@ 3.3 Feature extraction and selection
[…]
@@
4 Available AI Models for HEP Jet classification
Many DL architectures have been proposed in the SOTA of HEP domain to identify particles. Some of these architectures require input data in the form of images, while others utilize PC representations [37]. Table 5 summarizes and compares the most efficient ML and DL models, used in HEP, based on their architectures and performances.
[…]
ML, especially DL, has a rich historical presence in the field of particle physics. The concept of applying neural networks for tasks like distinguishing quarks and gluons, tagging Higgs particles, and identifying particle tracks has been around for more than two and a half decades. Nevertheless, the recent advancements in DL and the increased computational capabilities offered by graphics processing units have led to a significant enhancement in image recognition technology. As a result, there has been a renewed and heightened interest in utilizing these techniques. In the subsequent sections, we provide an overview of SOTA methods in both ML and DL. Figure 4 depicts a taxonomy of existing ML and DL techniques, summarizes the reviewed AI-based Jet classification models (discussed in Section 4), preprocessing and datasets (discussed in Section 3), and metrics (discussed in Section 2).
Figure 4: Taxonomy of ML and DL-based HEP techniques for Jet classification, with associated preprocessing, metrics, simulation tools and datasets.
4.1 ML-based methods
ML-based analysis of HEP Jet tagging has become an important technique in recent years. Jets are collimated sprays of particles, i.e, emitted from a source in a way that they are parallel or nearly parallel to each other, produced in high-energy particle collisions. Analyzing their properties is crucial for understanding the underlying physics processes. Jet images and PC are essentially 2D and 3D representations of the energy distribution within a Jet, where each pixel corresponds to a small region of the Jet. For example, Bogatskiy et al. in [38] introduced PELICAN, an ML architecture for particle physics that leveraged permutation-equivariant and Lorentz-invariant techniques, along with elementary equivariant aggregators and dense message-passing blocks. It processed 4-vector inputs representing particle jets as point clouds and employed a classifier to reduce rank-2 input arrays (pairwise dot products of 4-momentum vectors of particles in a jet) to permutation-invariant scalars using trace and total sum aggregation functions. Dense layers and a cross-entropy loss function were then used for optimization. Additionally, the PELICAN regressor predicted 4-momentum of particles using a permutation- and Lorentz-equivariant architecture with rank-preserving transformations and loss functions based on relative momentum and mass resolutions. Evaluation metrics included accuracy, AUC, background rejection rate, and relative resolutions. PELICAN achieved state-of-the-art performance in Jet classification, outperforming methods like LorentzNet while using approximately five times fewer parameters (45k only). Its low complexity, enhanced by equivariant aggregation, message-passing mechanisms, and its ability to handle regression tasks, made it suitable for real-time applications. However, its limitations included evaluation on limited datasets and reliance on hyperparameter tuning.
ML technique have been used in [39] by applying the shapley additive explanations (SHAP) method to explain the output of two HEP events ML classifiers (XGBoost and DNN) using the Higgs dataset. It demonstrates SHAP’s utility in understanding complex ML systems, particularly in the context of HEP event classifiers. The TreeExplainer and DeepExplainer methods from the Python SHAP library were used to compute SHAP values, revealing that features like $m\_bb$, $m\_wwbb$, and $m\_wbb$ were crucial in both models, although their distribution of SHAP values differed, indicating distinct learning processes. The process of extracting SHAP values are depicted in Figure 5.
In addition, quantum machine learning (QML) methods have recently found applications in addressing challenges within HEP, including separating signal from background [40], detecting anomalies [41], and reconstructing particle tracks [42].
Blance and Spannowsky [40] proposed a hybrid variational quantum classifier that combines quantum computing methods with classical neural network techniques to improve classification performance in particle physics research. The algorithm is applied to a resonance search in di-top final states, and it outperforms both classical neural networks and QML methods trained with non-quantum optimization methods. The classifier’s ability to be trained on small amounts of data indicates its potential benefits in data-driven classification problems. The proposed methodology was applied to the generated dataset, and the hybrid approach using the FST metric outperformed both classical neural networks and QML methods trained with non-quantum optimization methods in terms of maximizing learning outcomes; its accuracy can reach 72.6%. The hybrid approach also learned faster than an equivalent classical neural network or the classically trained variational quantum classifier. The paper [43] discusses the potential applications of quantum computation and QML in HEP, rather than focusing on deep mathematical structures. The authors claim that statistical ML methods are used for track and vertex reconstruction. These methods vary depending on the detector geometry and magnetic field used in the experiment. ML can help address these challenges by providing efficient and accurate methods for pattern recognition and particle identification. They suggest that quantum algorithms could potentially improve upon existing methods by offering faster and more efficient solutions to challenging problems in experimental HEP, such as particle identification and track reconstruction. This can be realized by creating a dataset recorded on tape through grid computing, which can be distributed for offline analysis using QML to extract information about particle trajectories developed inside the detectors. The work [44] investigates the potential of QML in HEP analysis at the LHC. The authors compare the performance of the quantum kernel algorithm to classical ML algorithms using 15 input variables and up to 50,000 events. They used 60 statistically independent datasets of 20,000 events each for their analysis. The AUC is used as the metric, and the results show that the performance of all methods improves with increasing dataset size. For 15 qubits, the quantum SVM-Kernel algorithm performs similarly to the classical support vector machine (SVM) and classical BDT algorithms. The quantum SVM-Kernel performances from the three different quantum computer simulators (Google, IBM, and Amazon) are comparable. The authors also claim that when a selection is implemented, permitting a signal acceptance rate of 70%, it results in the rejection of approximately 92% of background events, as indicated by the AUC. Consequently, the $S/\sqrt{B}$ ratio will experience an enhancement of approximately 150% compared to a scenario without any selection. Similarly, the researchers in [26] present a new approach to Jet classification using QML. The method involves embedding data into a quantum state, passing it through a variational quantum circuit, and performing a training procedure by minimizing a classical loss function. Probability measurements of the final state are then used to perform the classification. By exploiting the intrinsic properties of quantum computation, such as superposition and entanglement, the team aims to identify if a Jet contains a hadron formed by a $b$ or $\bar{b}$ quark at the moment of production. The approach could lead to new insights and enhance the classification performance in particle physics experiments. Two datasets have been used in this research: the complete dataset and the muon dataset, both of which belong to CERN. In the muon dataset analysis, 60,000 Jets are used for training and 40,000 Jets are used for testing. The muon dataset is a subset of the complete dataset, and it is used to evaluate the dependence of the quantum algorithms’ performance on the number of training events and the circuit complexity. The researchers compare the performance of their QML approach with that of DNN, long short-term memory (LSTM), and LSTM with convolutional layer models. They show that the results for tagging power as a function of the Jet $p_{T}$ and $\eta$ are comparable within the MSE error, and therefore, they consider only the DNN model for comparison with QML algorithms.
table[h!] A summary of available ML and DL architectures for Jet HEP classification. Ref. Year Model Input Performance TT Performance QG Other performance Link nature Acc. AUC Acc AUC Acc AUC [45] 2017 TopoDNN Image 0.916 0.972 No [46] 2018 CNN tagger Image 0.87 (DTJ) 0.943 (DTJ) No 0.945 (CJ) 0.988 (CJ) [47] 2019 PFN-ID PC 0.932 0.9819 0.9005 No [48] 2020 LGN PC 0.929 0.9640 0.803 0.8324 Yes [^9] https://github.com/fizisist/LorentzGroupNetwork [49] 2020 ParticleNet PC 0.940 0.9858 0.840 0.9116 No [50] 2021 EGNN PC 0.922 0.9760 0.803 0.8806 Yes [^10] https://github.com/vgsatorras/egnn [51] 2021 PCT PC 0.940 0.9855 0.841 0.9140 No [31] 2022 LorentzNet PC 0.942 0.9868 0.844 0.9156 No [52] 2022 PartT PC 0.944 0.9877 0.852 0.9230 Yes [^11] https://github.com/jet-universe/particle_transformer [38] 2022 PELICAN PC 0.942 0.9869 Yes [^12] https://github.com/abogatskiy/PELICAN [53] 2024 CGENNs PC 0.942 0.9869 Yes [^13] https://github.com/DavidRuhe/clifford-group-equivariant-neural-networks [54] 2024 L-GATr PC 0.942 0.9870 Yes [^14] https://github.com/Qualcomm-AI-research/geometric-algebra-transformer [55] 2024 MIParT-L PC 0.944 0.9878 0.853 0.9237 Yes [^15] https://github.com/jet-universe/particle_transformer
Abbreviation: TT: Top tagging; QG: Quark-gluon; DTJ: DeepTop Jets; CJ: CMS Jets; PC: Point cloud.
Table 5: A summary of available ML and DL architectures for Jet HEP classification, including columns for biases, generalizability, and recommended use cases. Bias levels range from moderate (limited datasets) to high (overfitting, dataset reliance), while generalizability is categorized as high (broad applicability), moderate (adequate performance with some limitations), and low (poor performance or untested on other tasks).
Ref. Year Model IN Acc. TT AUC TT Acc. QG AUC QG Acc. Other AUC Other Link Biases General. Recommended scenarios
[45] 2017 TopoDNN Image 0.916 0.972 No M L Top quark identification
[46] 2018 CNN tagger Image 0.87 (DTJ) 0.943 (DTJ) No H H Jet substructure
[47] 2019 PFN-ID PC 0.932 0.981 0.900 No L L Energy flow studies
[48] 2020 LGN PC 0.929 0.964 0.803 0.832 Yes [^16] https://github.com/fizisist/LorentzGroupNetwork L M Lorentz invariance studies
[49] 2020 ParticleNet PC 0.940 0.985 0.840 0.911 No M H Point cloud analysis
[50] 2021 EGNN PC 0.922 0.976 0.803 0.880 Yes [^17] https://github.com/vgsatorras/egnn L M Graph neural networks
[51] 2021 PCT PC 0.940 0.985 0.841 0.914 No L H Point cloud processing
[31] 2022 LorentzNet PC 0.942 0.986 0.844 0.915 No L M Lorentz group studies
[52] 2022 PartT PC 0.944 0.987 0.852 0.923 Yes [^18] https://github.com/jet-universe/particle_transformer L H Analysis of long-range feature dependencies in particles
[38] 2022 PELICAN PC 0.942 0.986 Yes [^19] https://github.com/abogatskiy/PELICAN M L Particle cloud matching
[53] 2024 CGENNs PC 0.942 0.986 Yes [^20] https://github.com/DavidRuhe/clifford-group-equivariant-neural-networks L H Clifford group analysis
[54] 2024 L-GATr PC 0.942 0.987 Yes [^21] https://github.com/Qualcomm-AI-research/geometric-algebra-transformer M H Geometric algebra studies
[55] 2024 MIParT-L PC 0.944 0.987 0.853 0.923 Yes [^22] https://github.com/jet-universe/particle_transformer L H Analysis of long-range feature dependencies in particles
Abbreviation: Input nature (IN); Point cloud (PC); Top tagging (TT); Quark-gluon (QG); DeepTop Jets (DTJ); CMS Jets (CJ); Moderate (M); High (H); Low (L).
4.2 MLP and DNN-based methods
[…]
MLP is an artificial neural network composed of multiple layers of nodes, including an input layer, one or more hidden layers, and an output layer. Each node in one layer is connected to every node in the subsequent layer. multi-Layer perceptron (MLP) can handle complex nonlinear relationships between input and output data, making them suitable for various tasks. MLPs are versatile, scalable, and can be trained using back-propagation, enabling them to learn from large datasets effectively and generalize well to unseen data. Kinematic parameters describe the motion of particles, including velocity, momentum ($p_{T,J}$) and trimmed Jet momentum ($p_{T,J,trim}$), energy, Jet mass $m_{J}$ and Jet mass trimmed $m_{J,trim}$ , and angles of emission, commonly used in physics and engineering analyses. Chakraborty et al. in [56] employed both kinematics and spectral function, which typically refers to a function that describes the distribution of energy or momentum states of particles in a particular physical system, to feed MLP classifier as described in Figure 6. The authors aim is to trim/discard Jet that are unlikely to have originated from the process of interest (effects of background noise). This selective removal helps to improve the accuracy of measurements and analyses by focusing only on the most relevant particles within a jet.